# On Mixed-Precision Iterative Methods and Analysis for Nearly Completely Decomposable Markov Processes

### v1

@@ 6.2 Empirical Performance Results
Figure 3: Performance comparison of our modified KMS algorithms (KMS+IR, KMS+Ri) and the KMS algorithm for diagonal matrices $\mathbf{P}_{ii}$ with dimension $n_{i}=200$ and number of blocks $m=20$, as a function of $\epsilon$ for the off-diagonal matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$. Performance results represent averages over $10$ independent trial runs.
@@ 7 Conclusions
We consider in this paper the problem of computing the stationary distribution of NCD Markov processes, a well-established area in the classical theory of Markov processes with broad applications in the design, modeling, analysis and optimization of computer systems. Our focus is on the design and analysis of numerical methods that address the performance bottlenecks of existing approaches and provide significant computational improvements to support the very large scale of current and emerging computer systems and applications as well as their potential deployment in online or real-time environments. In particular, we design general classes of algorithmic solution approaches that exploit forms of mixed-precision computation to significantly reduce computation times and exploit forms of iterative approximate methods to address the impact of inaccurate computations, further reduce computation times, and guarantee convergence. We then derive a mathematical analysis of our general classes of algorithmic approaches that rigorously establishes theoretical results on approximation errors, convergence behaviors, and other algorithmic properties. Numerical experiments empirically demonstrate that our general algorithmic approaches provide orders of magnitude improvements in the computation times to determine the stationary distribution of NCD Markov processes over the most efficient existing numerical methods which are based on aggregation and disaggregation.
@@ Appendix A Full Proofs of Some Theoretical Results
From the specific supposition $\bm{\pi}^{(t-1)}_{i}=\bm{\pi}_{i}+O(\varphi)$ together with $10<z\leq\|\bm{\pi}_{i}\|_{1}<$ 1, we obtain
@@
Given the theorem statements $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ and $\mathbf{s}=\mathbf{s}\mathbf{R}$, we take the difference between the former statement and the latter statement on both sides and derive
@@
where (a) follows upon substituting (17) into (18) and multiplying both sides on the right by $\mathbf{S}$. Given the theorem statement $\|\mathbf{s}\|_{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$ (according to the rightmost equation in (9)) and $\|\mathbf{S}\|=O(1)$ from the rightmost equation in (11), together with $\mathbf{R}^{(t-1)}=\mathbf{R}_{ij}+O(\varphi\epsilon)$ established above, we obtain
@@
It therefore follows from the last equation in (11) that $(\mathbf{s}^{(t-1)}-\mathbf{s})=O(\varphi)$, thus completing the proof. ∎
A.2 Proof of Theorem 4.2
Proof.
From (1), we have
@@
from which the desired result (16) follows.
Next, let the error in the approximate solution $\bm{\pi}^{(t-1)}$ at iteration $t-1$ of Algorithm 1 be of order $\varphi$, namely $\bm{\pi}^{(t-1)}=\bm{\pi}+O(\varphi)$. Then, from Lemma 4.1, we have
@@
It then follows from the initial arguments in the proof of Lemma 4.1 that
$\bm{\pi}^{(t)}=\bm{\pi}+O(\varphi\epsilon),\qquad\forall i\in[m],$
which yields the desired result. ∎

### v2

@@ 6.2 Empirical Performance Results
Figure 3: Performance comparison of our Algorithm 1 and Algorithm 2 with the KMS baseline for diagonal matrices $\mathbf{P}_{ii}$ having dimension $n_{i}=500$ and number of blocks $m=20$, as a function of $\epsilon\in\{0.01,0.05,0.1,0.15,0.2\}$ for the off-diagonal matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$. The comparative performance results represent averages over $10$ independent trial runs executed on a simulated H100100100100 GPU architecture.
Figure 3 presents the representative set of performance result comparisons for the case where the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, are fixed to be $500$, the number $m$ of diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, is fixed to be $20$, and the magnitude $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$, is varied such that $\epsilon\in\{0.01,0.05,0.1,0.15,0.2\}$. We first observe from Figure 3(a) that the mean number of iterations under Algorithm 1 are either identical to that of the baseline KMS algorithm or differ by at most one, with the mean number of iterations under Algorithm 2 somewhat higher requiring several additional iterations (though of lower computational complexity). The mean number of iterations as a function of the magnitude $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$, remains relatively flat for both KMS and Algorithm 1, as expected again due to the shared properties of the outer loop of all three algorithms. However, as observed in Figure 3(a), the mean number of iterations under Algorithm 2 increases and requires more outer-loop iterations as $\epsilon$ increases. Due to the nature of the block-Jacobi preconditioner, the convergence of the RI method becomes slower with increasing $\epsilon$ because now the part of the iteration matrix that is left outside the block-diagonal structure becomes more computationally significant. From Figure 3(b), we observe that both of the representative examples of our general algorithmic framework provide orders of magnitude reduction in the mean computation time with such performance improvements remaining relatively flat across the various magnitudes $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$. The mean computation time for each algorithm remains relatively flat as a function of the magnitude $\epsilon$, which is as expected due to the shared properties of the outer loop of all three algorithms. We note that our general approach in Algorithm 1 provides somewhat lower mean computations times than Algorithm 2 for larger values of $\epsilon$. This is due to the nature of the block-Jacobi preconditioner, as described above.
Next, we turn to present in Figure 4 the results from a representative sample of our second collection of numerical experiments based on selecting the $n_{i}\times n_{i}$ diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, from a standard library of matrices for real-world linear systems [33]. In particular, Figure 4 presents the representative set of performance result comparisons for the case where the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, consist of a particular real-world matrix in computational chemistry selected from the standard library with $n_{i}=730$, the number $m$ of diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, is fixed to be $20$, and the magnitude $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$, is varied such that $\epsilon\in\{0.01,0.05,0.1,0.15,0.2\}$. We first observe from Figure 4(a) that the mean number of iterations under Algorithm 1 differs from the baseline KMS algorithm by one, with the mean number of iterations under Algorithm 2 somewhat higher requiring several additional iterations (though of lower computational complexity). The mean number of iterations as a function of the magnitude $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$, remains flat for both KMS and Algorithm 1, as expected again due to the shared properties of the outer loop of all three algorithms. However, as observed in Figure 4(a), the mean number of iterations under Algorithm 2 increases and requires more outer-loop iterations as $\epsilon$ increases. Due to the nature of the block-Jacobi preconditioner, the convergence of the RI method becomes slower with increasing $\epsilon$ because now the part of the iteration matrix that is left outside the block-diagonal structure becomes more computationally significant. From Figure 4(b), we observe that both of the representative examples of our general algorithmic framework provide orders of magnitude reduction in the mean computation time with such performance improvements remaining relatively flat across the various magnitudes $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$, with some differences between Algorithm 1 and Algorithm 2 due to the block-Jacobi preconditioner reasons noted above. The mean computation times for each algorithm remains relatively flat as a function of the magnitude $\epsilon$, which is as expected due to the shared properties of the outer loop of all three algorithms. We note that our general approach in Algorithm 1 provides somewhat lower mean computations times than Algorithm 2 for larger values of $\epsilon$. This is again due to the nature of the block-Jacobi preconditioner, as described above. In particular, increasing $\epsilon$ makes the block off-diagonal couplings that are not captured by the preconditioner more significant.
(a) Mean number of iterations.
(b) Mean computation times (seconds).
Figure 4: Performance comparison of our Algorithm 1 and Algorithm 2 with the KMS baseline for diagonal matrices $\mathbf{P}_{ii}$ from real-world linear systems having dimensions $n_{i}=730$ and number of blocks $m=20$, as a function of $\epsilon\in\{0.01,0.05,0.1,0.15,0.2\}$ for the off-diagonal matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$. The comparative performance results represent averages over $10$ independent trial runs executed on a simulated H100100100100 GPU architecture.
@@ 7 Conclusions
We consider in this paper the problem of computing the stationary distribution of NCD Markov processes, a well-established area in the classical theory of Markov processes with broad applications in the design, modeling, analysis and optimization of computer systems and applications. Our focus is on the design and analysis of numerical methods that address the performance bottlenecks of the most-efficient existing approaches and provide significant computational improvements to support the very large scale of current and emerging computer systems and applications as well as their potential deployment in online or real-time environments. In particular, we devise a general mathematical framework of numerical solution methods that exploits forms of advances in mixed-precision computation to significantly reduce the performance bottlenecks associated with solving systems of linear equations and that exploits forms of advances in iterative approximate computing approaches to mitigate the impact of inaccurate computations, further reduce computation times, and guarantee convergence. We then derive a mathematical analysis that rigorously establishes theoretical properties of our general algorithmic framework including results on approximation errors, convergence behaviors, and other algorithmic characteristics. Numerical experiments based on validated simulation of mixed-precision computation empirically demonstrate that representative examples of our general algorithmic framework provide orders of magnitude improvements in the computation times to determine the stationary distribution of NCD Markov processes over the most-efficient existing numerical methods which are based on aggregation-disaggregation.
@@ A.1 Proof of Lemma 17
From the specific supposition $\bm{\pi}^{(t-1)}_{i}=\bm{\pi}_{i}+O(\varphi)$ together with $10<\tau\leq\|\bm{\pi}_{i}\|_{1}<$ 1, we obtain
@@
Given the theorem hypotheses $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ and $\mathbf{s}=\mathbf{s}\mathbf{R}$, where the latter follows from (9), we take the difference between the former equation and the latter equation on both sides and derive
@@
where (a) follows upon substituting (18) into (19) and multiplying both sides on the right by $\mathbf{S}$. Given the theorem hypothesis $\|\mathbf{s}\|_{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$ (according to the rightmost equation in (9)) and $\|\mathbf{S}\|=O(1)$ from the rightmost equation in (11), together with $\mathbf{R}^{(t-1)}=\mathbf{R}_{ij}+O(\varphi\epsilon)$ established above, we obtain
@@
It therefore follows from the last equation in (11) that $(\mathbf{s}^{(t-1)}-\mathbf{s})=O(\varphi)$, thus completing the proof of (16).
Lastly, from (1), we have
@@
from which the desired result (17) follows. ∎
A.2 Proof of Theorem 4.2
Proof.
Let the error in the approximate solution $\bm{\pi}^{(t-1)}$ at iteration $t-1$ of Algorithm 1 be of order $\varphi$, namely $\bm{\pi}^{(t-1)}=\bm{\pi}+O(\varphi)$. Then, from Lemma 17, we have
@@
It then follows from the initial arguments in the proof of Lemma 17 that
$\bm{\pi}^{(t)}_{i}=\bm{\pi}_{i}+O(\varphi\epsilon),\qquad\forall i\in[m],$
which yields the desired result.
Next, under the process employed for selecting the precision of computations in Step $\mathbf{5}$ and Step $\mathbf{3}$ of Algorithm 1 with respect to the condition number and the norm of the matrix $\mathbf{A}$ of the linear system, we know that the assumptions of the mixed-precision analysis of the IR method by Moler [26] are satisfied. It then follows from the convergence results of Moler [26] that, under these conditions, the IR method is guaranteed to converge linearly with an approximation error that decreases by a factor of $O(\kappa(\mathbf{A}))$ in each iteration, thus completing the proof. ∎